### v1

# A Study of the Generalized Naming Game and Bayesian Naming Game as Dynamical Systems

Abstract
We study the $\beta$-model ($\beta$-NG) and the Bayesian Naming Game (BNG) as dynamical systems. By applying linear stability analysis to the dynamical system associated with the $\beta$-model, we demonstrate the existence of a non-generic bifurcation with a bifurcation point $\beta_{c}=1/3$. As $\beta$ passes through $\beta_{c}$, the stability of isolated fixed points changes, giving rise to a one-dimensional manifold of fixed points. Notably, this attracting invariant manifold forms an arc of an ellipse. In the context of the BNG, we propose modeling the Bayesian learning probabilities $p_{A}$ and $p_{B}$ as logistic functions. This modeling approach allows us to establish the existence of fixed points without relying on the overly strong assumption that $p_{A}=p_{B}=p$, where $p$ is a constant.
I Introduction
The naming game is a multi-agent model that aims to simulate the spontaneous emergence of consensus among agents who interact pairwise according to two possible processes: agreement and word/name learning Castellano et al. [2009]; Blythe [2015]; Baronchelli [2016]; Chen and Lou [2019]; Marchetti et al. [2020a]. In its simplest realization, the agents are allow to learn only two possible names, denoted here by the symbols A and B, according to some basic rules (we refer the reader to Refs. Baronchelli [2016]; Marchetti et al. [2020a] for their definitions). Furthermore, if the underlying topology on which the agents may lie, is neglected altogether Castello et al. [2009]; Marchetti et al. [2020a], we call such a model as to the minimal naming game (MNG).
The MNG model can be thought as a particular case of two variants: the generalized naming game ($\beta$-NG) Baronchelli et al. [2007] and the Bayesian naming game (BNG) Marchetti et al. [2020b, a, 2021]. Both variants add human features otherwise absent in agents of MNG. In $\beta$-NG, a parameter $\beta$ is introduced to take into account the probability of acknowledged influence Friedkin [1990], thus affecting the agreement processes, while in BNG the name learning processes are now dictated by the time-dependent probabilities $p_{A}\equiv p_{A}\left(t\right)$ and $p_{B}\equiv p_{B}\left(t\right)$ necessary for generalizing the concept $\mathcal{C}$ associated to words A and B, respectively  Tenenbaum [1999]; Tenenbaum and Griffiths [2001]; Murphy [2012]. It is worth noting here that such probabilities are a direct consequence of Bayes rule Jeffreys [1939]. Therefore, according to the latter model the agents become Bayesian-rational individuals that in an uncertain world find their best action using the laws of probability de Finetti [1989]; Sanborn and Chater [2016]. That being said, the minimal naming game can be recovered from the above two variants by setting their characteristic parameters, i.e. $\beta$ and $p_{A},p_{B}$, to unity as illustrated in cartoon of Fig. 1.
The dynamics of the naming game is routinely investigated through computer simulations of multi-agent systems. The simulations allow for the computation of the observables of interest, e.g. the convergence time, the success rate, averaging over several hundred realizations Marchetti et al. [2020a]. However, the differential equations dictating the time-evolution for each model, can be also derived, assuming the mean-field approximation, thus offering an alternative approach to study some features of consensus dynamics. In fact, the differential equations of these models form two-dimensional (2D) dynamical systems whose study has some evident advantages. First, due to their low dimensionality, the time-evolution is much simpler when compared to that of naming game models where multiple names need to be taken into account Ma et al. [2023]. Second, one can exploit the methods provided by the dynamical systems theory Strogatz [2018]; Hirsch et al. [2013], such as bifurcation diagrams and phase portraits for autonomous systems, to infer whether or not the consensus emerges without the need to perform the numerical simulations, a task that becomes computationally expensive when the system size $N$, i.e., the number of agents, increases.
In this paper, we shall pursue the latter approach, thereby finding novel interesting dynamical features that, to our best knowledge, have been overlooked so far.
[…]
In Section III, we shall address the issue related to the lack of spontaneous symmetry breaking for the dynamical system associated with BNG. This issue arises when attempting to find the solution by numerically integrating the respective first-order differential system, assuming the initial condition $I_{0}=\left(0.5,0.5\right)$ and setting $p_{A}=p_{B}=p$ where $p$ is a constant, simultaneously.
Denoting the solution as $\left(x,y\right)$, where $x\equiv x\left(t\right)$ and $y\equiv y\left(t\right)$ are the population fractions having either A or B in their inventories at time $t$ respectively, it is found that $\left(x,y\right)$ asymptotically converges to $\left(0.38,0.38\right)$. Thus, there is no convergence to consensus, similar to what occurs with MNG for the very same setup Castellano et al. [2009]; Marchetti et al. [2020a].
This is in contrast to what is found in multi-agent simulations, where consensus always emerges either at A or at B for MNG and BNG models. It has already been noted that such a discrepancy between simulations and dynamical systems arises from the fact that the dynamical system approach cannot account for the ubiquitous density fluctuations, or equivalently noise, present in numerical simulations Castellano et al. [2009]. However, this issue is particularly significant here, as the above setup was used to prove the existence of consensus for the dynamical system associated with BNG Marchetti et al. [2020b].
To address this serious problem, we propose modeling the Bayesian probabilities as time-dependent logistic functions (see Eq. 12James et al. [2023]. Subsequently, by numerically integrating the modified equations, we demonstrate that consensus consistently emerges at either A or B. This outcome is attributed to the symmetry breaking induced by the time-dependent probabilities, disrupting the invariance under the swap of variables $x$ and $y$ (see Eqs. 10 and  11). Accordingly, the solution closely resembles that obtained from multi-agent simulations with system size $N\gtrsim 200$ Marchetti et al. [2020b].
Figure 1: The cartoon depicts the relationship between the MNG, $\beta$-NG, and BNG through their characteristic parameters: $\beta\in[0,1]$ and the learning probabilities $p_{A},p_{B}\in[0,1]$. Setting $\beta$ and $p_{A},p_{B}$ to unity in $\beta$-NG and BNG models, respectively, one recovers the MNG.
II The $\beta$-NG model as Dynamical System
In mean-field approximation the time-evolution of $\beta$-NG model is dictated by the following system of first-order differential equations:
$\dot{x}=-xy+\beta\left(1-x-y\right)^{2}+Kx\left(1-x-y\right)\,,$ (1)
$\dot{y}=-xy+\beta\left(1-x-y\right)^{2}+Ky\left(1-x-y\right)\,,$ (2)
where $\dot{x},\dot{y}$ the derivatives of $x,y$, with respect to time. We set $K\equiv\left(3\beta-1\right)/2$ in Eqs. 1, 2 while $\beta\in[0,1]$. Next, setting $\dot{x}=0$ and $\dot{y}=0$ into Eqs.1 and 2, it was previously determined that three isolated fixed points exist: A, B, and CBaronchelli et al. [2007]; Castellano et al. [2009]. Applying the LST revealed a first-order non-equilibrium transition at $\beta=\beta_{c}$. Furthermore, the system converges to point C for $\beta<\beta_{c}$, leaving a fraction of undecided agents equal to $1-2c_{\beta}$, while a spontaneous emergence of consensus occurs at either A or B for $\beta>\beta_{c}$. These findings are summarized in the first and third columns of Table 1.
While the above analysis is certainly correct and has been confirmed by multi-agent simulations, it does not address the generic bifurcation that characterizes the dynamical system in question.
In the following, we shall unravel the dynamical features of the dynamical system by applying the LSA to the problem at hand in a systematic way. First, let us define the functions $f_{1}\left(x,y\right),f_{2}\left(x,y\right)$ such that $\dot{x}=f_{1}\left(x,y\right),\dot{y}=f_{2}\left(x,y\right)$ according to Eqs. 1, 2. It is then possible to compute the eiggenvalues of the following Jacobian matrix $J$ evaluated at a given fixed equilibrium point with Cartesian coordinates $\left(\tilde{x},\tilde{y}\right)$ Strogatz [2018]
$J=\begin{pmatrix}\partial_{x}f_{1}\left(\tilde{x},\tilde{y}\right)&\partial_{y}f_{1}\left(\tilde{x},\tilde{y}\right)\\ \partial_{x}f_{2}\left(\tilde{x},\tilde{y}\right)&\partial_{y}f_{2}\left(\tilde{x},\tilde{y}\right)\end{pmatrix}\,,$ (3)
where the symbols $\partial_{x}$, $\partial_{y}$ denote the partial derivatives of $f_{1}\left(x,y\right)$, $f_{2}\left(x,y\right)$ with respect to $x,y$, respectively.
Second, the local equilibrium character of each fixed point can be determined by the eigenvalues of the corresponding matrix $J$. In particular, their equilibrium character is easily identified by looking at the inequalities between the eigenvalues as shown in the trace-determinant plane (also called the bifurcation diagram) (see Fig. 2)  Hirsch et al. [2013]; Perez-Garcia et al. [2019]).
Figure 2: The types of the equilibria for the dynamical system with Eqs. 1, 2 and $\beta\in[0,1]$ according to LSA, in the trace-determinant plane. The respective inequalities between the two (real) eigenvalues $\lambda_{i},\lambda_{j}$ ($i\neq j$) of the matrix $J$ evaluated at each type of fixed point are also shown. The parabola in the (black) solid line corresponds to $\operatorname{tr}\left(J\right)^{2}-4\det\left(J\right)=0$. Note that all the saddle points lie on quadrants III and IV of the trace-determinant plane.
The matrix $J$’s entries evaluated at the fixed point A read: $J_{11}=-K$, $J_{12}=-1-K$, $J_{21}=0$, $J_{22}=-1$ while the matrix entries takes the values $J_{11}=-1$, $J_{12}=0$, $J_{21}=-1-K$, $J_{22}=-K$ when computed at B. Since in both cases the traces and the determinants of the Jacobian matrices are the same, i.e. $\operatorname{tr}\left(J\right)=-1-K$ and $\det\left(J\right)=K$, one must conclude that these matrices have same eigenvalues. Therefore, the fixed points $A,B$ present the very same stability behaviour for each value of $\beta$ in $[0,1]$. Accordingly, in the following we shall focus on the equilibrium character of A only, denoting its corresponding eigenvalues by the symbols $\lambda_{1},\lambda_{2}$. By means of the equations $\operatorname{tr}\left(J\right)=\lambda_{1}+\lambda_{2}$ and $\det\left(J\right)=\lambda_{1}\lambda_{2}$, one immediately finds these two real eigenvalues: $\lambda_{1}=-1$ and $\lambda_{2}=-K$. Since $\lambda_{1}$ is always negative, both eigenvalues are negative when $\beta>\beta_{c}$, while $\lambda_{2}>0$ when $\beta<\beta_{c}$ (see Fig. 3). Looking at the classification scheme in Fig. 2, one must conclude that $A,B$ are saddle points and asymptotically stable nodes for $\beta\in[0,\beta_{c})$ and $\beta\in(\beta_{c},1]$, respectively. So far, the results agree with those reported in Ref. Baronchelli et al. [2007]. However, at $\beta=\beta_{c}$, $\lambda_{2}=0$, and hence both A and B must belong to the stable line corresponding to the negative half-axis (see the (orange) solid line in Fig. 2). Therefore, A and B cannot be longer considered isolated, but according to LSA there must exist an infinite number of other fixed points. We will return to this point later.
Figure 3: The real eigenvalues $\lambda_{i}$ with $i=1,2$ and $i=3,4$ corresponding to the Jacobian matrix J (Eq. 9) evaluated at A and $C$, respectively, as functions of the parameter $\beta$. The vertical (dotted) line corresponds to $\beta=\beta_{c}$.
We now turn our attention to the equilibrium character of the fixed point $C$, denoting its corresponding eigenvalues by $\lambda_{3}$ and $\lambda_{4}$. To this end, we shall compute them by means of the following formula  Sagués and Epstein [2003]
$\lambda_{3},\lambda_{4}=\frac{1}{2}\left(\operatorname{tr}\left(J\right)\pm\sqrt{\operatorname{tr}\left(J\right)^{2}-4\det\left(J\right)}\right)\,.$ (4)
In Eq. 4 we plot $\lambda_{3},\lambda_{4}$ as function of $\beta$ by means of Eq.. 4. From a direct comparison of the corresponding curves with those of $\lambda_{1},\lambda_{2}$, one immediately concludes that $C$ has the opposite equilibrium character of A and B for each each $\beta\in[0,1]\setminus\{\beta_{c}\}$, see Fig. 3. However, when $\beta=\beta_{c}$, $C$ belongs to the stable line as well because $\lambda_{3}=0$ while $\lambda_{4}$ negative (see Fig. 2).
Noting that the stable line is characterized by $\operatorname{tr}\left(J\right)<0$ and $\det\left(J\right)=0$, the last important result could also be obtained without the need to compute the eigenvalues, as follows. First, we note that at $C=\left(c_{\beta},c_{\beta}\right)$, the matrix $J$’s entries must satisfy the identities $J_{11}=J_{22}$ and $J_{12}=J_{21}$, thus the Jacobian matrix is now symmetric. Second, $J_{11},J_{12}$ take the following form as functions of $\beta$
$J_{11}=-\left(\frac{\beta+1}{2}\right)+\left(\frac{1-\beta}{2}\right)c_{\rm\beta}\,,$ (5)
and
$J_{12}=-2\beta+\left(\frac{5\beta-1}{2}\right)c_{\rm\beta}\,,$ (6)
respectively.
Therefore, in such a case $\operatorname{tr}\left(J\right)=2J_{11}$ and $\det\left(J\right)=J_{11}^{2}-J_{12}^{2}$. Now noting that $\det\left(J\right)=\left(J_{11}+J_{12}\right)\left(J_{11}-J_{12}\right)$, one finds that
$J_{11}-J_{12}=K\left(1-2c_{\beta}\right)\,.$ (7)
So, as $K=0$ when $\beta=\beta_{c}$, it follows that the left hand side of Eq. 7, and hence $\det\left(J\right)$ are both zero. Next, we note that the corresponding trace is negative, i.e. $\operatorname{tr}\left(J\right)=2/3\left(c_{\beta}-2\right)<0$, because $c_{\beta}=2-\sqrt{3}\approx 0.26$.
Overall, our analysis shows that the fixed points $A,B,C$ exist for all values of the parameter $\beta$ but there an exchange of stabilities between $A,B$ and $C$ as $\beta$ passes through $\beta_{c}$. In this regard, we can say that the dynamical system undergoes a generic bifurcation similar to that of the 1D differential equation discussed in Section I. For the sake of completeness, we summarize all these findings in Table 1.
Table 1: The classification of the fixed points $A,B,C$ according to the linear stability theory. Note that when the isolated equilibrium points are stable, they are also asymptotically stable. The fixed points $A,B,C$ are no longer isolated at bifurcation point $\beta=\beta_{c}$.
Fixed Point $\beta\in[0,\beta_{c})$ $\beta=\beta_{c}$ $\beta\in(\beta_{c},1]$
A saddle point stable line stable node
B saddle point stable line stable node
C stable node stable line saddle point
Next, the existence of a continuum of fixed points at the bifurcation point could be easily inferred by examining the streamline plot of the velocity field $\mathbf{v}=\left(\dot{x},\dot{y}\right)$ corresponding to Eqs. 1, 2. To this end, in Fig. 4 we plot such a velocity field on the model’s domain in Cartesian $xy$-plane: $\{(x,y)\in\mathbb{R}^{2}:0\leq x\leq 1,0\leq y\leq 1,0\leq x+y\leq 1\}$. It is then evident that there exists one-dimensional manifold of fixed points (orange solid line), and that all trajectories approach it along parallel lines. This finding is consistent with the stable line, that is, the negative negative half-axis of trace $\operatorname{tr}J$, obtained from LSA (see Fig. 2).
Figure 4: Streamline plot of $\mathbf{v}=\left(\dot{x},\dot{y}\right)$ corresponding to Eqs. 1, 2 at bifurcation point $\beta=\beta_{c}$. Here the stable line (orange solid line) corresponds to the graph of the map $x\mapsto 1/2\left(-\sqrt{3}\sqrt{4x-x^{2}}+x+2\right)$ with $x\in[0,1]$. The important initial condition $I_{0}=\left(0.5,0.5\right)$ is also shown.
We now prove that the above one-dimensional manifold constitutes an arc of an ellipse. First, its Cartesian equation can be obtained by setting $\beta=\beta_{c}$ and either $\dot{x}=0$ or $\dot{y}=0$ into Eqs. 1, 2. One then finds that the set of points must satisfy the following second-order degree polynomial equation in the variables $x,y$
$x^{2}-xy+y^{2}-2x-2y+1=0\,.$ (8)
According to Eq. 8 the one-dimensional manifold corresponds to the graph of the function $x\mapsto 1/2\left(-\sqrt{3}\sqrt{4x-x^{2}}+x+2\right)$ with $x\in[0,1]$. Next, we construct the symmetric matrix $M$ of the quadratic form associated with Eq. 8 in order to classify it as a nondegenerate conic section Apostol [1969]. In the present case, one finds
$M=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}\,,$ (9)
with determinant $\det\left(M\right)=3/4$. As a consequence its eigenvalues must the same sign, so Eq. 8 yield a conic section that is an ellipse with semi-major axis, semi-minor axis and eccentricity corresponding to $a=\sqrt{6}$, $b=\sqrt{2}$ and $e=\sqrt{2/3}$, respectively. The graph of such an ellipse with its center $0=\left(2,2\right)$ and foci $F_{1}=\left(2-\sqrt{2},2-\sqrt{2}\right)$ and $F_{2}=\left(2+\sqrt{2},2+\sqrt{2}\right)$ is shown in Fig. 5. Note that in such a case, point $C$ now has coordinates $\left(c_{\beta_{c}},c_{\beta_{c}}\right)$ with $c_{\beta_{c}}=2-\sqrt{3}$ and hence lies on the ellipse between points A and B.
Figure 5: The graph of the ellipse with Eq. 8 emerging at bifurcation point $\beta=\beta_{c}$. The ellipse’s center and foci are $0=\left(2,2\right)$, $F_{1}=\left(2-\sqrt{2},2-\sqrt{2}\right)$ and $F_{2}=\left(2+\sqrt{2},2+\sqrt{2}\right)$, respectively. The fixed points $A=\left(1,0\right)$, $B=\left(0,1\right)$ and $C=\left(c_{\beta_{c}},c_{\beta_{c}}\right)$ where $c_{\beta_{c}}=2-\sqrt{3}$ are no longer isolated equilibria.
We conclude this section by noting another important consequence of the existence of a non-generic bifurcation. The dynamical system approach provides a deterministic mean-field description of the $\beta$-model, but the trajectories from its multi-agent simulations are intrinsically noisy. Therefore, one would expect that there is some effect of noise close to the bifurcation Erban et al. [2009].
To illustrate this, let us consider $x$, or equivalently $y$, obtained from a single run of a simulation starting from $I_{0}$ as a random variable $X_{t}$. In Fig. 6, we plot six randomly chosen sample paths of $X_{t}$ as functions of time from simulations with $N=1,000$ agents. Note that the paths labeled by $\omega=1,2,3$ and $\omega=4,5,6$ correspond to $\beta=0.2$ and $\beta=\beta_{c}$, respectively where each $\omega_{i}$ could be understood as a specific outcome in the sample space.
The paths with $\omega=1,2,3$ exhibit a pattern completely different from those at the bifurcation point, i.e., with $\omega=4,5,6$. While the former remain close to the sample mean $E[X_{t}]$ (black solid line) computed from $1,200$ trajectories, thus showing the typical mean-returning behavior of the Ornstein-Uhlenbeck (OU) process Uhlenbeck and Ornstein [1930]; Vasicek [1977]; Oksendal [2010], the latter seem to wander in a more random fashion. Preliminary statistical analysis Marchetti [prep] confirms that the sample paths at $\beta=0.2$ belong to OU process, while those at bifurcation point to a nonstandard Brownian motion Einstein [1905]; Langevin [1908]; Wiener [1923]; Ritort [2019].
Figure 6: Randomly chosen sample paths of $X_{t}$ ($X_{0}=0.5$) as functions of time $t$ from multi-agents simulations with $N=1,000$ agents. The paths with label $\omega=1,2,3$ and $\omega=4,5,6$ correspond to $\beta=0.2$ and $\beta=\beta_{c}$, respectively. Solid lines in black color and red color corresponds to sample mean $E[X_{t}]$ computed from $1,200$ trajectories for $\beta=0.2$ and $\beta=\beta_{c}$, respectively.
III The Bayesian Naming Game as a Dynamical System
According to the BNG model, agents must learn new, unknown names through their previous experiences and background knowledge. Furthermore, since the learning process occurs under uncertainty, a probabilistic approach is necessary for modeling such a process. In this context, a Bayesian framework proposed by Tenenbaum and co-workers seems appropriate Tenenbaum [1999]; Murphy [2012]; Griffiths and Tenenbaum [2006]; Griffiths and Kalish [2007]; Tenenbaum et al. [2011]; Perfors et al. [2011]; Lake et al. [2015]. As a result of this probabilistic framework, agents attempt to generalize by computing the learning probabilities $p_{A}$ and $p_{B}$, which depend on the positive examples stored by the agents during their pairwise interactions. In this regard, the BNG model paves the way for incorporating many human cognitive biases Tversky and Kahneman [1974]; Hahn [2014]; Sanborn and Chater [2016]; Ngampruetikorn and Stephens [2016]; Madsen et al. [2018] into the consensus dynamics, which were out of reach for the previous variants of naming game.
Assuming the mean-field approximation, the time evolution of BNG is governed by the following non-autonomous first-order differential system Marchetti et al. [2020b]:
$\dot{x}=-p_{B}xy+\left(1-x-y\right)^{2}+\frac{3-p_{B}}{2}x\left(1-x-y\right)\,,$ (10)
$\dot{y}=-p_{A}xy+\left(1-x-y\right)^{2}+\frac{3-p_{A}}{2}y\left(1-x-y\right)\,.$ (11)
Note that finding a solution to the above system is not possible without making some assumptions about the probabilities $p_{A}\left(t\right)$ and $p_{B}\left(t\right)$, both falling within the interval [0,1] with an unknown functional form. In Ref.Marchetti et al. [2020b], it was proposed to set $p_{A}=p_{B}=p$, thus making the probabilities time-independent. This assumption, along with the initial condition $I_{0}$, played a crucial role in estimating an upper bound of about $20\%$ for the number of agents with two names in their inventories Marchetti et al. [2021]  [^2] The agents having two names in their inventories are often called bilinguals in literature.. However, it cannot be used at all to demonstrate the existence of the fixed points A and B in full generality, as wrongly assumed in Ref.Marchetti et al. [2020b]. In fact, consensus cannot be reached in such a setup due to the absence of symmetry breaking, as previously discussed within the MNG in Section I.
That being said, we cannot infer the exact time-dependent probability pattern of $p_{A},p_{B}$ directly from the simulations of BNG, because it is strongly affected by the system size $N$. However, there is useful information in them, independent of $N$, that can be advantageously exploited. Indeed, it is found that $p_{A},p_{B}$ grow monotonically with time as the success rate observable Marchetti et al. [2020b]. Furthermore, name learning is a few-shot process, where agents successfully generalize from a small number of positive examples Tenenbaum [1999]; Marchetti et al. [2021]. Therefore, the functional form of $p_{A}\left(t\right)$ and $p_{B}\left(t\right)$ cannot be too smooth. With the above considerations in mind, we propose modeling both $p_{A}$ and $p_{B}$ as time-dependent logistic function (or sigmoid function) Murphy [2012]. The logistic function has a S-shaped form that can be modified appropriately thanks to its two characteristic parameters. Denoting it by the symbol $g$, this function reads:
$g\left(t\right)=\frac{1}{1+\exp\left(-\eta\left(t-t_{m}\right)\right)}\,,$ (12)
where $\eta$ is the logistic growth rate (or the steepness of the curve), and $t_{m}$ corresponds to the value of $t$ of function’s midpoint.
Next, we have to tune the parameters of the logistic functions to ensure that the Bayesian probabilities yield the right solution for the problem at hand. In this regard, we need to examine how the Bayesian probabilities behave in multi-agent simulations, assuming that at the beginning $50\%$ of the agents know the word A and the remainder know the other word B Marchetti et al. [2020b]  [^3] From the dynamical system point of view this setup corresponds exactly to have the initial condition $I_{0}$.. In Fig. 7, we plot the cumulative distribution functions (CDF) from the values of $p_{A}$ and $p_{B}$ obtained from a single run of a system with size $N=3,000$ with the above initial condition. Examining these curves derived from simulation data, it is evident that the probabilities $p_{A}$ and $p_{B}$ cannot be assumed to be equal throughout the entire time-evolution of the system, but it may be possible only for a short time interval (see discussion below).
Furthermore, we note that what diversifies the cumulative distribution function (CDF) is the initial bias introduced in the simulations. In order to distinguish the synonyms A and B, different thresholds, $n^{\ast}{ex,A}=5$ and $n^{\ast}{ex,B}=6$, of the number of positive examples stored by the agents, necessary for computing the probabilities $p_{A}$ and $p_{B}$, respectively, are assigned. As a result, this bias will favor the learning of name A. There is also a step-like progression toward unity in the cumulative distribution functions as a direct consequence of few-shot learning Tenenbaum [1999].
Figure 7: Cumulative distribution functions for the values of $p_{A},p_{B}$ obtained from a simulation of the system with size $N=3,000$, starting from the initial condition $I_{0}$ and with Erlang prior. Other simulation parameters are $n^{\ast}_{ex,A}=5$, $n^{\ast}_{ex,B}=6$. Due to this setup, the consensus is reached at A.
That being said, we need to tune the parameters of two logistic functions, here denoted by $g_{A}$ and $g_{B}$, in order to differentiate the time-dependence of probabilities $p_{A}$ and $p_{B}$, respectively. For example, we can set the same value $\eta=2$ for both logistic functions, but at least the parameter $t_{m}$ should have different values to distinguish $g_{A}$ from $g_{B}$. To this end, we set $t_{m,A}=10$ and $t_{m,B}=130$ for $g_{A}$ and $g_{B}$, respectively (see inset of Fig. 8). This parameter choice ensures a step-like pattern for the probabilities, making them non-smooth, while also favoring the learning of name A. With these parameters in place, we expect that the time-evolution of the dynamical system reproduces the main features observed in simulations of multi-agent system.
To confirm the above conjecture we proceed to numerically integrate Eqs. 10, 11, replacing $p_{A}$ with $g_{A}$ and $p_{B}$ with $g_{B}$. The numerical result shows that the solution $\left(x,y\right)$ asymptotically converge to $A=\left(1,0\right)$ (see Fig. 8444We attempted to numerically integrate the differential system using both the Runge-Kutta method and the LSODA algorithm. We found that both methods produced identical solution.. Furthermore, by swapping the values of parameters $t_{m,A}$ and $t_{m,B}$, it is also straightforward to show that there exist the consensus at $B=\left(0,1\right)$. Overall, there exist a spontaneous emergence of consensus despite the dynamics follows a mean-field deterministic prescription.
Remarkably, the numerical integration also yields a solution $\left(x,y\right)$ that closely resembles that obtained from a simulation with relatively large number of agents, i.e., $N\gtrsim 200$ (see Ref. Marchetti et al. [2020b]).
We conclude noting that the solution $\left(x,y\right)$ splits in its two components $x$ and $y$ after a transient time during which $x=y$, see Fig. 8. Such an initial stage of dynamics, where $x=y$, can be thought as a dynamical system whose differential equations are: $\dot{x}=-p_{B}xy$, $\dot{y}=-p_{A}xy$. This fact clearly explains that the dynamics of BNG dynamical system is primarily dictated by the Bayesian probabilities $p_{A}$ and $p_{B}$, at least in the initial stage of the dynamics. Once again, it is their interplay and time-dependence that lead to the subsequent symmetry-breaking process, essential for the emergence of consensus.
Figure 8: Solution of Eqs. 10, 11 assuming $g_{A}$ and $g_{B}$ with parameters $\eta=2$, $t_{m,A}=10$ and $t_{m,B}=130$, The differential equations at hand were numerically integrated by means of LSODA algorithm assuming $I_{0}=\left(0.5,0.5\right)$.
Acknowledgments
The author would like to express gratitude to Steven Strogatz for his helpful insights on the bifurcation and to Konstantinos “Kostas” Zygalakis for the important remarks about the effect of noise around the critical points and bifurcations and Andrea Baronchelli for reading the manuscript. Some results presented here were obtained during the Spring of 2021 while the author was at the National Institute of Chemical Physics and Biophysics, Tallinn, Estonia. Therefore, the author extends thanks to Martti Raidal for the financial support via the European Regional Development Fund CoE program TK133 “The Dark Side of the Universe”, and to Els Heinsalu and Marco Patriarca for proposing the naming game as a research topic.